OpenStudy (anonymous):

Write the equation of the line with a slope of 2 that passes through the points (—2, —7) and (5, 7).

OpenStudy (anonymous):

@mathstudent55

OpenStudy (anonymous):

I know y=mx+b. Any help?

OpenStudy (anonymous):

@Data_LG2 This is my last question

OpenStudy (anonymous):

i think it's y = 2x + 3

OpenStudy (anonymous):

The choices are: y – 7 = 2x + 5 y = 2x y = 2x + 3 y = 2x – 3

OpenStudy (anonymous):

let's if you're right... \(\sf (Y-Y_1)=m(X-X_1)\) Find m=2 plug the points into the equation: \(\sf (Y-7)=2(X-5)\\Y=2x-10+7\\Y=2x-3\) confirm using the other point: \(\sf (Y+7)=2(X+2)\\Y=2x+4-7\\Y=2x-3\)

OpenStudy (anonymous):

It's 'minus 3' thanks for all your help! @Data_LG2

OpenStudy (anonymous):

no problem (: